# Bone Soups: A Seek-and-Soup Model Merging Approach for Controllable Multi-Objective Generation

### v1

@@ Abstract
User information needs are often highly diverse and varied. A key challenge in current research is how to achieve controllable multi-objective generation while enabling rapid adaptation to accommodate diverse user demands during test time. Existing solutions, such as Rewarded Soup, focus on merging language models individually tuned on single objectives. While easy to implement and widely used, these approaches face limitations in achieving optimal performance due to their disregard for the impacts of competing objectives on model tuning. To address this issue, we propose Bone Soup, a novel model merging approach that first seeks a series of backbone models by considering the impacts of multiple objectives and then makes the soup (i.e., merge the backbone models). Specifically, Bone Soup begins by training multiple backbone models for different objectives using multi-objective reinforcement learning. Each backbone model is guided by a combination of backbone reward signals. To ensure that these models are optimal for the Pareto front, the backbone rewards are crafted by combining standard reward functions into basis vectors, which can then be modified through a rule-based construction method. Bone Soup leverages a symmetric circulant matrix mapping to generate the merging coefficients, which are used to merge the backbone models according to user preferences. Extensive experimental results demonstrate that Bone Soup exhibits strong controllability and Pareto optimality in controllable multi-objective generation, providing a more effective and efficient approach to addressing diverse user needs at test time.
@@
Guofu Xie1, Xiao Zhang1, Ting Yao2, Yunsheng Shi2, 1Gaoling School of Artificial Intelligence, Renmin University of China 2Tencent {guofuxie, zhangx89}@ruc.edu.cn
1 Introduciton
Human preferences and their information needs are highly diverse, and even for the same task, users may have distinct personalized demands in different scenarios Wu et al. (2024); Rame et al. (2023); Wang et al. (2024a); Shi et al. (2024); Chen et al. (2024). This diversity introduces a significant controllability challenge for AI service providers, who must develop machine learning models that can effectively adapt to a wide range of individual preferences. A key area of research addressing this challenge is controllable multi-objective generation (CMOG), which focuses on guiding the behavior of language models (LMs) to meet diverse and real-time user requirements without the need for retraining Zhang et al. (2023); Rame et al. (2023); Shi et al. (2024); Wang et al. (2024a). For example, Bing Copilot offered users modes like “More Accurate”, “More Balanced”, and “More Creative”, allowing for customization based on their discrete requirements.
@@ 1 Introduciton
A straightforward approach to implementing CMOG is through prompt-based control Dong et al. (2023); Ramnath et al. (2023); Yang et al. (2024); Wang et al. (2024a), where LMs are guided to generate content according to user preferences for different objectives by modifying only the input prompts. These approaches can be seen as an implicit control mechanism since it does not modify the model parameters at test time. Recently, some explicit approaches of controlling LMs have gained attention, known as model merging Wortsman et al. (2022); Rame et al. (2023); Tang et al. (2024); Yu et al. (2024); Yadav et al. (2024); Yang et al. (2023); Wang et al. (2024b); Ilharco et al. (2022). In model merging approaches, model parameters from different LMs are combined at test time to accommodate varying user preferences. This form of test-time adaptation often provides more reliable control for CMOG, as it achieves control at the parameter level.
@@ 2.2 Problem Analyses
We also provide a comparison of the disparity between solutions of different methods and the oracle in Figure 3. Therefore, constructing appropriate backbone rewards to train the backbone models is crucial for achieving Pareto optimality and controllability in CMOG.
@@ 3.1 Approach Overview
We design and implement a more sophisticated merging-based approach Bone Soup for CMOG. Instead of directly interpolating between original base models, we propose to first seek the backbone models which ensures better Pareto optimality, and then determine the merging coefficients to contribute to better controllability. Figure 2 illustrates the overall workflow of our method.
@@ 3.2.1 Obtaining Backbone Models
As mentioned earlier, introducing additional reward models may help restructure better backbone models, we introduce several rules to efficiently and effectively determine matrix $\bm{B}$ :
@@ 4.1 Experiments Setups
Implementation Details. We use LLama-2 7B Touvron et al. (2023) for Helpful Assistant task and Reddit Summary task and use T5-large Raffel et al. (2020) for Long Form QA task. For all three tasks, we choose the best $\beta\in\{0.8,0.7,0.6\}$ by only training for 20% total steps and evaluate the hypervolume. As for the extrapolation of $\hat{\bm{\theta}}$, we also select the optimal $\alpha\in\{0.1,0.2,0.3,0.4,0.5\}$ using a validation datasets following the approach in Zheng et al. (2024).
Baselines. We consider three most typical latest CMOG approaches including prompt-based approach Rewards-in-Context (RiC) Yang et al. (2024), decoding-time approach MOD Shi et al. (2024) and merging-based method Rame et al. (2023) and follow their settings of the Hyperparameters. Detailed introduction and discussion about baselines are in Appendix A.3.1.
@@ 4.2.1 Long Form QA task
In Figure 4 and Table 2, we compare Bone Soup (BS) at different $\beta$ values with Rewarded Soup (RS) and MORLHF. The selection of $\beta$ is discussed in A.2.6 and A.2.2. BS consistently outperforms RS and closely approximates Oracle MORLHF across three trade-offs. In factuality vs completeness and relevance vs completeness, BS even surpasses MORLHF, achieving a superior Pareto front.
Additionally, experiments in A.2.1 show that combining multiple rewards generally improves the backbone model. This led us to investigate merging backbone models based on user preferences, without considering the reward interplay in merging coefficients. As seen in Figure 4, the direct merge approach ("ABA") performs worse than RS in factuality vs completeness, but slightly outperforms it in the other two trade-offs, though still behind BS.
@@ A.1.2 Model Merging
In the field of multi-task learning, various model merging approaches have been proposed. Tang et al. (2024) using a dynamic routing mechanism trained in test-time adaptation to determine the model fusion coefficients. Some approaches exploit model parameter redundancy, leading to pruning-based approaches Yadav et al. (2024); Yu et al. (2024). AdaMerging Yang et al. (2023) employs an unsupervised approach to train merging coefficients and adjusts the merging coefficients at different layers of the models. TALL-masks Wang et al. (2024b) generates a task-specific mask matrix using a predefined threshold derived from independent models. The key distinction between our approach and the above works lies in that they are not designed for, nor capable of, achieving controllable generation. In contrast, we have developed a series of techniques specifically aimed at optimizing for Pareto optimality and controllability.
@@
A.2 Additional Experiments
A.2.1 RQ1: The Significance of Reconstructing Appropriate Backbone Models.
@@ A.2 Additional Experiments
The effectiveness of model merging fundamentally depends on the quality and diversity of the backbone models. Relying solely on models trained for specific objectives may not yield optimal results, as such models might not fully explore or exploit the entire reward landscape. In Figure 4, we observe that the specializing models extrapolated by the two backbone models of Bone Soup consistently extend in two reward dimensions and outperform the specializing models tuned in RS, verifying the fact that models tuned with a specific reward may not always be the optimal ones for that reward and can be outperformed by models derived through various interpolation techniques.
@@
A.2.2 RQ2: Can the Small-Scale Selection Yield a $\beta$ Consistent with or Near the Optimal $\beta$?
@@
A.2.3 RQ3: Can Our Bone Soup Matrix Construction Method Effectively Identify High-Quality Matrices in a Vast Solution Space?
@@
A.2.4 RQ4: Robustness Analysis of Bone Soup
@@
A.2.5 RQ5: How Does the Determination of Merging Coefficients Affect the Performance of Bone Soup?

### v2

@@ Abstract
User information needs are often highly diverse and varied. A key challenge in current research is how to achieve controllable multi-objective generation while enabling rapid adaptation to accommodate diverse user demands during test time. Existing solutions, such as Rewarded Soup, focus on merging language models individually tuned on single objectives. While easy to implement and widely used, these approaches face limitations in achieving optimal performance due to their disregard for the impacts of competing objectives on model tuning. To address this issue, we propose Bone Soup, a novel model merging approach that first seeks a series of backbone models by considering the impacts of multiple objectives and then makes the soup (i.e., merge the backbone models). Specifically, Bone Soup begins by training multiple backbone models for different objectives using multi-objective reinforcement learning. Each backbone model is guided by a combination of backbone reward signals. To ensure that these models are optimal for the Pareto front, the backbone rewards are crafted by combining standard reward functions into basis vectors, which can then be modified through a rule-based construction method. Bone Soup leverages a symmetric circulant matrix mapping to generate the merging coefficients, which are used to merge the backbone models according to user preferences. Extensive experimental results demonstrate that Bone Soup exhibits strong controllability and Pareto optimality in controllable multi-objective generation, providing a more effective and efficient approach to addressing diverse user needs at test time. Code is available at https://github.com/andyclsr/BoneSoups.
@@
Guofu Xie1, Xiao Zhang1thanks: Corresponding author: Xiao Zhang., Ting Yao2, Yunsheng Shi2 1Gaoling School of Artificial Intelligence, Renmin University of China, Beijing, China 2Tencent {guofuxie, zhangx89}@ruc.edu.cn{tessieyao, yunshengshi}@tencent.com
1 Introduction
@@
Human preferences and their information needs are highly diverse, and even for the same task, users may have distinct personalized demands in different scenarios Wu et al. (2024); Rame et al. (2023); Wang et al. (2024a); Shi et al. (2024); Chen et al. (2024). This diversity introduces a significant controllability challenge for AI service providers Shen et al. (2024a); Chen et al. (2023); Shen et al. (2024b), who must develop learning models that can effectively adapt to a wide range of user preferences. A key area of research addressing this challenge is controllable multi-objective generation (CMOG), which focuses on guiding the behavior of language models (LMs) to meet diverse and real-time user requirements without the need for retraining Zhang et al. (2023); Rame et al. (2023); Shi et al. (2024); Wang et al. (2024a).
For example, Bing Copilot offered users modes like “More Accurate”, “More Balanced”, and “More Creative”, allowing for customization based on their discrete requirements.
A straightforward approach to implementing CMOG is through prompt-based control Dong et al. (2023); Ramnath et al. (2023); Yang et al. (2024); Wang et al. (2024a), where LMs are guided to generate content according to user preferences for different objectives by modifying only the input prompts. These approaches can be seen as an implicit control mechanism since it does not modify the model parameters at test time. Recently, some explicit approaches of controlling LMs have gained attention, known as model merging Wortsman et al. (2022); Rame et al. (2023); Tang et al. (2024); Yu et al. (2024); Yadav et al. (2024); Yang et al. (2023); Wang et al. (2024b); Ilharco et al. (2022). In model merging approaches, model parameters from different LMs are combined at test time to accommodate varying user preferences.
This form of test-time adaptation often provides more reliable control for CMOG, as it achieves control at the parameter level.
@@ 2.2 Problem Analyses
We also provide a comparison of the disparity between solutions of different methods and the oracle in Figure 3.
Through the above example, we have demonstrated that BoneSoup can achieve solutions closer to the oracle. To further illustrate this point, we present the following theorem, which provides a lower bound on the interval where BoneSoup outperforms Rewarded Soup. This result proves that the front obtained by BoneSoup is, in most cases, superior to that of Rewarded Soup. We follow the setting in Rame et al. (2023) using quadratic reward functions and with Hessians proportional to identity matrices to derive the theorem.
Theorem 1.
Given quadratic reward functions with Hessians proportional to identity matrices:
$r_{i}(\theta)=r_{i}(\theta_{i})-k_{i}\|\theta-\theta_{i}\|^{2},i\in\{1,2\},$
where $k_{i}\in\mathbb{R}_{+}$ are distinct,and $\theta_{i}$ is the global maximum for reward $r_{i}$. Let the reward combination weight matrix be $B=\begin{pmatrix}\beta&1-\beta\\ 1-\beta&\beta\end{pmatrix},\beta\in(\frac{1}{2},1)$, then the backbone rewards of the bone-soup approach can be denoted as $[h_{1},h_{2}]^{T}=B[r_{1},r_{2}]^{T}$.Let $\bm{\mu}=[\mu,1-\mu]^{T}$ be the user preference and the testing reward can written as $g_{\bm{\mu}}(\theta):=\bm{\mu}^{T}\begin{bmatrix}r_{1}\\ r_{2}\end{bmatrix}$.Denote the approximate solutions for the testing reward $g_{\mu}(\theta)$ of the soup-like approach and the bone-soup approach as $\bar{\theta}$ and $\bar{\theta}^{bone}$, respectively.Then,for any fixed $\beta\in(\frac{1}{2},1)$, when $\mu\in\left(\frac{1-\sqrt{2\beta^{2}-2\beta+1}}{2},\frac{1+\sqrt{2\beta^{2}-2\beta+1}}{2}\right)$,
$g_{\mu}(\bar{\theta})<g_{\mu}(\bar{\theta}^{bone}),$
with interval length $\sqrt{2\beta^{2}-2\beta+1}\geq\frac{\sqrt{2}}{2}$.
Therefore, constructing appropriate backbone rewards to train the backbone models is crucial for achieving Pareto optimality and controllability in CMOG.
Proof.
Please refer to Appendix A.2. ∎
@@ 3.1 Approach Overview
We design and implement a more sophisticated merging-based approach Bone Soup for CMOG. Instead of directly interpolating between original base models, we propose to first seek the backbone models which ensures better Pareto optimality, and then determine the merging coefficients to contribute to better controllability.
Figure 2 illustrates the overall workflow of our method.
@@ 3.2.1 Obtaining Backbone Models
As mentioned earlier, introducing additional reward models may help restructure better backbone models, we introduce several rules
to efficiently and effectively determine matrix $\bm{B}$ :
@@ 4.1 Experiments Setups
Implementation Details. We use LLama-2 7B Touvron et al. (2023) for Helpful Assistant task and Reddit Summary task and use T5-large Raffel et al. (2020) for Long Form QA task.
For all three tasks, we choose the best $\beta\in\{0.8,0.7,0.6\}$ by only training for 20% total steps and evaluate the hypervolume.
As for the extrapolation of $\hat{\bm{\theta}}$, we also select the optimal $\alpha\in\{0.1,0.2,0.3,0.4,0.5\}$ using a validation datasets following the approach in Zheng et al. (2024).
Baselines. We consider three latest CMOG approaches including prompt-based approach Rewards-in-Context (RiC) Yang et al. (2024), decoding-time approach MOD Shi et al. (2024) and merging-based method Rame et al. (2023) and follow their settings of the Hyperparameters. Detailed introduction and discussion about baselines are in Appendix A.4.1.
@@ 4.2.1 Long Form QA task
In Figure 4 and Table 2, we compare Bone Soup (BS) at different $\beta$ values with Rewarded Soup (RS) and MORLHF. The selection of $\beta$ is discussed in A.3.6 and A.3.2. BS consistently outperforms RS and closely approximates Oracle MORLHF across three trade-offs. In factuality vs completeness and relevance vs completeness, BS even surpasses MORLHF, achieving a superior Pareto front.
Additionally, experiments in A.3.1 show that combining multiple rewards generally improves the backbone model. This led us to investigate merging backbone models based on user preferences, without considering the reward interplay in merging coefficients. As seen in Figure 4, the direct merge approach ("ABA") performs worse than RS in factuality vs completeness, but slightly outperforms it in the other two trade-offs, though still behind BS.
@@ 7 Limitations
8 Acknowledgement
This work was partially supported by the National Natural Science Foundation of China (No. 62376275, 92470205). Work partially done at Beijing Key Laboratory of Research on Large Models and Intelligent Governance, and Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE. Supported by fund for building world-class universities (disciplines) of Renmin University of China.
@@ A.1.2 Model Merging
In the field of multi-task learning, various model merging approaches have been proposed. Tang et al. (2024) using a dynamic routing mechanism trained in test-time adaptation to determine the model fusion coefficients. Some approaches exploit model parameter redundancy, leading to pruning-based approaches Yadav et al. (2024); Yu et al. (2024).
AdaMerging Yang et al. (2023) employs an unsupervised approach to train merging coefficients and adjusts the merging coefficients at different layers of the models.
TALL-masks Wang et al. (2024b) generates a task-specific mask matrix using a predefined threshold derived from independent models.
The key distinction between our approach and the above works lies in that they are not designed for, nor capable of, achieving controllable generation.
In contrast, we have developed a series of techniques specifically aimed at optimizing for Pareto optimality and controllability.
@@
A.2 The Proof of Theorem 1
Theorem 1.
Given quadratic reward functions with Hessians proportional to identity matrices:
$r_{i}(\theta)=r_{i}(\theta_{i})-k_{i}\|\theta-\theta_{i}\|^{2},i\in\{1,2\},$
where $k_{i}\in\mathbb{R}_{+}$ are distinct,and $\theta_{i}$ is the global maximum for reward $r_{i}$. Let the reward combination weight matrix be $B=\begin{pmatrix}\beta&1-\beta\\ 1-\beta&\beta\end{pmatrix},\beta\in(\frac{1}{2},1)$, then the backbone rewards of the bone-soup approach can be denoted as $[h_{1},h_{2}]^{T}=B[r_{1},r_{2}]^{T}$.Let $\bm{\mu}=[\mu,1-\mu]^{T}$ be the user preference and the testing reward can written as $g_{\bm{\mu}}(\theta):=\bm{\mu}^{T}\begin{bmatrix}r_{1}\\ r_{2}\end{bmatrix}$.Denote the approximate solutions for the testing reward $g_{\mu}(\theta)$ of the soup-like approach and the bone-soup approach as $\bar{\theta}$ and $\bar{\theta}^{bone}$, respectively.Then,for any fixed $\beta\in(\frac{1}{2},1)$, when $\mu\in\left(\frac{1-\sqrt{2\beta^{2}-2\beta+1}}{2},\frac{1+\sqrt{2\beta^{2}-2\beta+1}}{2}\right)$,
$g_{\mu}(\bar{\theta})<g_{\mu}(\bar{\theta}^{bone}),$
with interval length $\sqrt{2\beta^{2}-2\beta+1}\geq\frac{\sqrt{2}}{2}$.
Proof.
The testing reward $g_{\mu}=\mu r_{1}+(1-\mu)r_{2}$ is quadratic thus has an unique global maximum $\theta^{*}$,that we find analytically:
$\nabla_{\theta}g_{\mu}(\theta)=0$ $\Rightarrow\mu k_{1}(\theta-\theta_{1})+(1-\mu)k_{2}(\theta-\theta_{2})=0$
$\Rightarrow\theta^{*}=\frac{\mu k_{1}\theta_{1}+(1-\mu)k_{2}\theta_{2}}{\mu k_{1}+(1-\mu)k_{2}}$
The approximate solution $\bar{\theta}$ for the testing reward $g_{\mu}$ of the soup-like approach is formulated as,
$\bar{\theta}=\mu\theta_{1}+(1-\mu)\theta_{2}$
Consider the bone-soup approach,we have the backbone rewards and their corresponding global maximums as follows, $h_{1}=\beta r_{1}+(1-\beta)r_{2},\theta_{1}^{bone}=\frac{\beta k_{1}\theta_{1}+(1-\beta)k_{2}\theta_{2}}{\beta k_{1}+(1-\beta)k_{2}}$ $h_{2}=(1-\beta)r_{1}+\beta r_{2},\theta_{2}^{bone}=\frac{(1-\beta)k_{1}\theta_{1}+\beta k_{2}\theta_{2}}{(1-\beta)k_{1}+\beta k_{2}}$
the merging coefficients can be calculated as $\bm{\lambda}=B^{-1}\bm{\mu}=[\lambda,1-\lambda]^{T}$,where $\lambda=\frac{\beta+\mu-1}{2\beta-1}$,then we have the soup-like approach’s approximate solution $\bar{\theta}^{bone}$ for the testing reward $g_{\mu}$ as,
$\bar{\theta}^{bone}=\lambda\theta_{1}^{bone}+(1-\lambda)\theta_{2}^{bone}$
We set the error function as $E(\beta,\mu)=\|\bar{\theta}^{bone}-\theta^{*}\|^{2}$.Since the soup-like approach can be regarded as a special case of the bone-soup approach when $\beta=1$,we use $E(1,\mu)$ to denote the error of the soup-like approach.Under the current settings, the testing reward $g_{\mu}$ can be written as $g_{\mu}=c_{1}-c_{2}\|\theta-\theta^{*}\|^{2}$,where $c_{1}$ and $c_{2}$ are constants, and $c_{2}\in\mathbb{R}_{+}$.Therefore $g_{\mu}(\bar{\theta})<g_{\mu}(\bar{\theta}^{bone})\Leftrightarrow E(\beta,\mu)<E(1,\mu)$.
The expressions for $E(\beta,\mu)$ and $E(1,\mu)$ can be calculated as follows,
$E(\beta,\mu)=$
$\|\lambda\frac{\beta k_{1}\theta_{1}+(1-\beta)k_{2}\theta_{2}}{\beta k_{1}+(1-\beta)k_{2}}+$
$(1-\lambda)\frac{(1-\beta)k_{1}\theta_{1}+\beta k_{2}\theta_{2}}{(1-\beta)k_{1}+\beta k_{2}}-\frac{\mu k_{1}\theta_{1}+(1-\mu)k_{2}\theta_{2}}{\mu k_{1}+(1-\mu)k_{2}}\|^{2}$
$=$ $\left(\frac{k_{1}k_{2}(k_{1}-k_{2})(\beta-\mu)(\beta+\mu-1)}{(\mu k_{1}+(1-\mu)k_{2})(\beta k_{1}+(1-\beta)k_{2})((1-\beta)k_{1}+\beta k_{2})}\right)^{2}\|\theta_{1}-\theta_{2}\|^{2}$
$E(1,\mu)=(\frac{(k_{1}-k_{2})(1-\mu)\mu}{\mu k_{1}+(1-\mu)k_{2}})^{2}\|\theta_{1}-\theta_{2}\|^{2}$
[…]
$E(\beta,\mu)<E(1,\mu)$
$\Leftrightarrow(\frac{k_{1}k_{2}(\beta-\mu)(\beta+\mu-1)}{[\beta k_{1}+(1-\beta)k_{2}][(1-\beta)k_{1}+\beta k_{2}]})^{2}<((1-\mu)\mu)^{2}$
since
$~{}[\beta k_{1}+(1-\beta)k_{2}][(1-\beta)k_{1}+\beta k_{2}]$
$=k_{1}k_{2}[2\beta^{2}-2\beta+1+\beta(1-\beta)(\frac{k_{1}}{k_{2}}+\frac{k_{2}}{k_{1}})]$
$\geq k_{1}k_{2}(2\beta^{2}-2\beta+1+2\beta(1-\beta))=k_{1}k_{2}$
Therefore, we obtain:
$\left(\frac{k_{1}k_{2}(\beta-\mu)(\beta+\mu-1)}{[\beta k_{1}+(1-\beta)k_{2}][(1-\beta)k_{1}+\beta k_{2}]}\right)^{2}$ $<\left((\beta-\mu)(\beta+\mu-1)\right)^{2}$
$E(\beta,\mu)<E(1,\mu)$
$\Leftarrow(\mu(1-\mu))^{2}-((\beta-\mu)(\beta+\mu-1))^{2}$
$=(\beta-\beta^{2})(-2\mu^{2}+2\mu+\beta^{2}-\beta)>0$
We can observe that,for any fixed $\beta\in(\frac{1}{2},1)$, the right-hand side of the equation holds, for all $\mu\in\left(\frac{1-\sqrt{2\beta^{2}-2\beta+1}}{2},\frac{1+\sqrt{2\beta^{2}-2\beta+1}}{2}\right)$, i.e. $E(\beta,\mu)<E(1,\mu)$ holds.Besides,we can find that the interval length $\sqrt{2\beta^{2}-2\beta+1}\geq\frac{\sqrt{2}}{2}$.Thus, the theorem is proved.
A.3 Additional Experiments
A.3.1 RQ1: The Significance of Reconstructing Appropriate Backbone Models.
@@ A.3 Additional Experiments
The effectiveness of model merging fundamentally depends on the quality and diversity of the backbone models. Relying solely on models trained for specific objectives may not yield optimal results, as such models might not fully explore or exploit the entire reward landscape.
In Figure 4, we observe that the specializing models extrapolated by the two backbone models of Bone Soup consistently extend in two reward dimensions and outperform the specializing models tuned in RS, verifying the fact that models tuned with a specific reward may not always be the optimal ones for that reward and can be outperformed by models derived through various interpolation techniques.
@@
A.3.2 RQ2: Can the Small-Scale Selection Yield a $\beta$ Consistent with or Near the Optimal $\beta$?
@@
A.3.3 RQ3: Can Our Bone Soup Matrix Construction Method Effectively Identify High-Quality Matrices in a Vast Solution Space?
@@
A.3.4 RQ4: Robustness Analysis of Bone Soup
@@
A.3.5 RQ5: How Does the Determination of Merging Coefficients Affect the Performance of Bone Soup?